# Insulating and metallic phases in the one-dimensional Hubbard-Su-Schrieffer-Heeger model: Insights from a backflow-inspired variational wave function

### v2

@@ I Introduction
The electron-phonon interaction plays an important role in condensed-matter physics, as it often leads to stark modifications of the low-energy electronic properties, the most striking example being the insurgence of superconductivity in a variety of materials de Gennes (1966). In addition, it may favor lattice deformations, turning, e.g., a metallic system into a (Peierls) insulator Peierls (1955). In molecular solids, the electron-phonon coupling is at the origin of the Jahn-Teller effect, which can lead to orbital order and geometric distortions Khomskii (2014). In high-temperature (cuprate) superconductors, the interplay between electron pairing and charge-density waves (the so-called stripes) has been widely discussed in the past 30 years and represents a crucial aspect that must be elucidated in order to reach a definitive understanding of those materials Kivelson et al. (2003); Fradkin et al. (2015).
@@
From a computational perspective, an ab-initio technique capable of accurately treating both electron-electron (e-e) and electron-phonon (e-ph) interactions is currently lacking, especially when effects beyond the Born-Oppenheimer approximation are relevant. In this respect, the definition of suitable low-energy models on a lattice may serve as a valuable framework for shedding light on the phenomena that emerge in this kind of problems. The single-band Hubbard model (or its strong-coupling version, the $t-J$ model) serves as the simplest example to capture electron correlations Arovas et al. (2022); Qin et al. (2022). Within this approach, phonons can be included by adding harmonic oscillators on each lattice site, leading to acoustic and/or optical branches. Then, the e-ph coupling can be modeled by Fröhlich, Holstein, or Su-Schrieffer-Heeger (SSH) Hamiltonians Fröhlich (1954); Holstein (1959); Su et al. (1979, 1980); Hohenadler and Fehske (2018). The former two cases are suitable to describe polaron effects in dielectric crystals; the latter one has been introduced to characterize solitons in one-dimensional materials (such as Polyacetylene). Here, lattice vibrations are directly coupled to the electron hopping.
@@
In one spatial dimension, whenever the electron band is half filled and in the absence of e-e interaction, the SSH model gives rise to a Peierls instability – a spontaneous dimerization of the chain, with alternating long and short bonds between nearest-neighbor sites – as obtained by earlier quantum Monte Carlo calculations and renormalization-group arguments Fradkin and Hirsch (1983). In the presence of an e-e interaction, such as the onsite Hubbard repulsion, the situation is more complicated. In the adiabatic limit (i.e., when taking an infinite ion mass and, therefore, no ion dynamics), lattice distortions appear for infinitesimal e-ph coupling Baeriswyl and Maki (1985), as also previously found in the strong-coupling (Heisenberg) model Cross and Fisher (1979). A full quantum mechanical treatment of the model, including both electron and ion dynamics, has been addressed in a relatively limited number of works. In Ref. Hirsch (1983), it has been argued that an on-site electron repulsion suppresses the Peierls distortions for large phonon frequencies. Later, a few works, mainly using Monte Carlo or density-matrix renormalization group (DMRG) techniques, analyzed a variety of models with both quantum lattice fluctuations and short-range e-e interactions Sengupta et al. (2003); Pearson et al. (2011); Weber et al. (2015). In particular, acoustic and optical phonons have been compared at half-filling (also considering different ways to couple lattice vibrations to the electronic hopping) Malkaruge Costa et al. (2023). The general outcome suggests that two phases can be stabilized: the Mott insulator (when the e-e interaction dominates), with no lattice distortions and gapless spin excitations, and the Peierls insulator (when the e-ph coupling dominates), characterized by lattice distortions, a two-fold degenerate ground state, and a fully gapped spectrum. The transition between them is described by the Kosterliz-Thouless universality class Giamarchi (2003), making its precise location extremely challenging.
@@
Away from half-filling, a limited number of investigations have been attempted. Indeed, ground-state Monte Carlo techniques suffer from the sign problem, which prevents one from assessing large system sizes, and the DMRG accuracy is strongly affected by the large entanglement of the gapless (metallic) ground state. One interesting aspect is to understand whether a so-called Luther-Emery phase Luther and Emery (1974), namely a metal with gapped spin excitations, may ever emerge when doping the Peierls insulator. In this respect, analytic and numerical works demonstrated that the Luther-Emery liquid appears in the doped two-leg ladder Hubbard model (with no phonons) Fabrizio (1996); Daul and Noack (1998); Nishimoto et al. (2008); Balents and Fisher (1996); Shen et al. (2023). Its existence is triggered by the presence of multiple Fermi points in the non-interacting band structure, as discussed in weak-coupling approaches Fabrizio (1996); Balents and Fisher (1996). For example, in the one-dimensional Hubbard model with both nearest- ($t$) and next-nearest-neighbor ($t^{\prime}$) hopping, a metallic phase with gapped spin excitations emerges by lightly doping the dimerized insulator found at $t^{\prime}>t/2$ Fabrizio (1996); Daul and Noack (1998); Nishimoto et al. (2008). By analogy, one could expect that the same phase could appear upon doping the (dimerized) Peierls insulator stabilized by the electron-phonon interaction. An early work Yonemitsu and Imada (1996), based upon bosonization and renormalization-group methods, indeed showed that a spin-gapped metal emerges close to half-filling due to the phonon-assisted backward scattering. More recently, both static and dynamical properties have been addressed by DMRG Banerjee et al. (2023). However, a single electron density $n$ has been considered, relatively far away from half-filling (i.e., $n=0.75$), and Luttinger-liquid properties are observed, with gapless charge and spin degrees of freedom.
@@
In this work, we study the one-dimensional Hubbard-SSH model by means of a variational Monte Carlo approach (VMC). We employ a variational wave function composed of three terms: the first one is a simple phonon condensate, which can give rise to a staggered (dimerized) pattern of the lattice displacements; the second one is a Bardeen-Cooper-Schrieffer (BCS) state; the last one is a Jastrow factor, which includes electron-electron, phonon-phonon, and electron-phonon correlations. The main novelty of the variational Ansatz resides in the inclusion of “backflow” correlations within the BCS wave function, through an auxiliary BCS Hamiltonian which explicitly depends upon the phonon coordinates. We present a detailed analysis of the SSH model in presence of a local Hubbard interaction, both at half-filling and for various hole doping. At half-filling, the ground state is either a Mott or a Peierls insulator, depending on the values of the e-e and e-ph couplings, in agreement with previous calculations Sengupta et al. (2003); Pearson et al. (2011); Weber et al. (2015). The central part of our work concerns the doped case, where metallic states appear. In fact, while doping a Mott insulator gives rise to a standard Luttinger liquid, similarly to what happens in the Hubbard model without phonons, in a lightly doped Peierls insulator the spin gap remains finite, thus stabilising a Luther-Emery liquid, which has never been noticed in this kind of coupled electron-phonon problems.
@@ II.2 The variational wave function
The Metropolis move includes either single or double (spin flip) electron hopping at nearest-neighbor sites and local updates of the displacements $X_{j}\mapsto X_{j}+\Delta$, with $\Delta$ uniformly distributed within the interval $[-\Delta_{\rm max},\Delta_{\rm max}]$.
Let us now discuss the specific construction of the variational wave function that is used in this work. In previous works Watanabe et al. (2015); Karakuzu et al. (2017); Ferrari et al. (2024), the electronic part has been taken as a pair-product state:
@@
where the hopping ($t_{ij}=t^{*}_{ji}$) and pairing ($\Delta_{ij}=\Delta_{ji}$) amplitudes are variational parameters, as well as the chemical potential $\mu$. Alternatively, it is possible to optimize directly the pairing function $f_{i,j}$, without passing through the definition of the auxiliary Hamiltonian Ohgoe and Imada (2017). Within these approaches, the electronic wave function does not depend upon the phonon displacements, e.g., the pairing function does not change along the Markov chain, once the variational parameters are kept fixed.
@@
where $g_{m}=g_{-m}$ and $h_{m}=-h_{-m}$ are variational parameters for $m=\pm 1$ and ±$\pm 3$ (suitable for a possible bond dimerization) and $\{X_{j}\}$ are the phonon displacements in the configuration $|X\rangle$ visited along the Markov chain. In this way, $\langle X;n_{\sigma}|\Psi_{\rm e}\rangle$ depends upon the actual phonon configuration. In this sense, a sort of “backflow” correlations are included in the electronic part of the variational wave function. In the standard case on the continuum, the effective position of every electron (from which the Slater determinant is constructed) depends on all the other ones Feynman and Cohen (1956); Lee et al. (1981); Schmidt et al. (1981); on the lattice, this approach has been extended by constructing single-particle orbitals or pairing functions that explicitly depend upon the many-body electron configuration Tocchio et al. (2008, 2011). Here, instead, backflow correlations involve electrons and phonons.
The phonon part has the same form as used in Ref. Ferrari et al. (2024):
@@
Finally, the e-e and e-ph correlations are included by standard Jastrow factors Ohgoe and Imada (2017); Ferrari et al. (2024):
@@
where the pseudo-potentials $v_{ij}$, $u_{ij}$, and $w_{ij}$ are treated as variational parameters. It turns out that the optimal energy is obtained by imposing both translation and reflection symmetries, with $v_{ji}=v_{ij}$, $u_{ji}=u_{ij}$, and $w_{ji}=-w_{ij}$. The density-density Jastrow factor is particularly relevant to describe the Mott insulating phase, in which the Fourier transform of the pseudo-potential $v_{q}\approx 1/q^{2}$ for small momenta $q$ Capello et al. (2005).
In the following, we consider clusters with $L$ sites and periodic-boundary conditions. All the variational parameters of the variational wave function are optimized by using the stochastic reconfiguration approach Sorella (2005); Becca and Sorella (2017). Our variational wave function is implemented in a JAX-based code Bradbury et al. (2018) that runs on parallel CPUs thanks to mpi4jax Häfner and Vicentini (2021).
@@ III.1 Benchmarks with DMRG calculations
First, we would like to assess the accuracy of our approach and compare the variational energies to DMRG results at half-filling, for $L=50$. In the DMRG calculations, periodic-boundary conditions are taken on the Hamiltonian, but a truncation in the phonon Hilbert space is necessary. Here, we fix a maximum occupancy of $10$ bosons per site (which is always much larger than the average occupation). Then, the energy of the electron-phonon Hamiltonian is optimized for a matrix-product state (which, for computational efficiency, has no translational symmetry) by means of the DMRG algorithm implemented in the ITensor library Fishman et al. (2022). The accuracy of DMRG calculations is verified by evaluating the variance of the total energy, which is always below $0.007t^{2}$.
@@ III.2 Results at half-filling
Let us now discuss the results at half-filling by focusing on the best variational Ansatz with backflow correlations. By varying the e-e repulsion $U$ and the e-ph interaction $\lambda$, the ground state is either a Mott (uniform) or a Peierls (dimerized) insulator Sengupta et al. (2003). Within our approach, the presence of a finite lattice distortion is signaled by the stabilization of a finite $z$ parameter in Eq. (13), which gives a dimerized pattern around which the phonon displacements are distributed. In this case, the translational symmetry is explicitly broken in the variational wave function and dimerization can be also detected from the calculation of the bond-order parameter:
@@
The results are summarized in the phase diagram shown in Fig. 2, fixing $\hbar\omega/t=1$. Here, the trivial limits are $\lambda=0$, which corresponds to the Mott insulator with no e-ph coupling, and $U=0$, which corresponds to the Peierls insulator. The transition between these two phases is of the Kosterliz-Thouless type Giamarchi (2003). Therefore, it is extremely difficult to locate its actual location within numerical calculations, since the spin gap of the Peierls phase is exponentially small close to the transition. This is particularly relevant in the vicinity of the non-interacting limit $U=\lambda=0$, where large clusters become necessary to detect the presence of the tiny dimerization that exists for $\lambda\to 0$. Nevertheless, sufficiently away from the transition line (and the non-interacting limit), our variational approach can straightforwardly distinguish two different regimes, with and without bond order.
@@
The results for $U/t=4$ are reported in Fig. 3 for two different regimes, i.e., $\lambda=0.04$ (Mott insulator) and $\lambda=0.25$ (Peierls insulator). The DMRG results (on the same cluster size, optimized for the Hamiltonian with periodic-boundary conditions) are also reported for comparison. First of all, the difference between Mott and Peierls states appears in the bond structure factor, which shows a huge peak at $q=\pi$ (i.e., $B(\pi)\approx 30$) in the latter case; the presence of a true dimerization implies a divergence of $B(\pi)$ with $L$. Within the variational approach, the huge peak at $q=\pi$ comes from the stabilization of a finite phonon parameter $z$ in Eq. (13), i.e. a finite lattice distortion. In other words, the existence of a Peierls phase can be more easily detected by looking at $z$ [or equivalently at $B_{\rm e}$ defined in Eq. (17)] than considering the bond structure factor $B(q)$. Within the Mott state, instead, $B(\pi)$ is much reduced, still having a cusp. Let us now turn to density and spin correlations. For both insulating phases, $N(q)$ is smooth and shows a quadratic behavior ($N(q)\propto q^{2}$) for small $q$, indicating that the charge excitations are gapped Capello et al. (2005). Also the spin structure factor $S(q)$ in the Peierls phase is smooth, signaling that also spin excitations are gapped. By contrast, in the Mott phase $S(q)$ is linear for small $q$ and has a peak at $q=\pi$, which is expected to diverge logarithmically with the system size and signals the presence of gapless spin excitations.

### v3

@@ I Introduction
The electron-phonon interaction plays an important role in condensed-matter physics, as it often leads to stark modifications of the low-energy electronic properties, the most striking example being the insurgence of superconductivity in a variety of materials [1]. In addition, it may favor lattice deformations, turning, e.g., a metallic system into a (Peierls) insulator [2]. In molecular solids, the electron-phonon coupling is at the origin of the Jahn-Teller effect, which can lead to orbital order and geometric distortions [3]. In high-temperature (cuprate) superconductors, the interplay between electron pairing and charge-density waves (the so-called stripes) has been widely discussed in the past 30 years and represents a crucial aspect that must be elucidated in order to reach a definitive understanding of those materials [4, 5].
@@
From a computational perspective, an ab-initio technique capable of accurately treating both electron-electron (e-e) and electron-phonon (e-ph) interactions is currently lacking, especially when effects beyond the Born-Oppenheimer approximation are relevant. In this respect, the definition of suitable low-energy models on a lattice may serve as a valuable framework for shedding light on the phenomena that emerge in this kind of problems. The single-band Hubbard model (or its strong-coupling version, the $t-J$ model) serves as the simplest example to capture electron correlations [6, 7]. Within this approach, phonons can be included by adding harmonic oscillators on each lattice site, leading to acoustic and/or optical branches. Then, the e-ph coupling can be modeled by Fröhlich, Holstein, or Su-Schrieffer-Heeger (SSH) Hamiltonians [8, 9, 10, 11, 12]. The former two cases are suitable to describe polaron effects in dielectric crystals; the latter one has been introduced to characterize solitons in one-dimensional materials (such as Polyacetylene). Here, lattice vibrations are directly coupled to the electron hopping.
@@
In one spatial dimension, whenever the electron band is half filled and in the absence of e-e interaction, the SSH model gives rise to a Peierls instability – a spontaneous dimerization of the chain, with alternating long and short bonds between nearest-neighbor sites – as obtained by earlier quantum Monte Carlo calculations and renormalization-group arguments [13]. In the presence of an e-e interaction, such as the onsite Hubbard repulsion, the situation is more complicated. In the adiabatic limit (i.e., when taking an infinite ion mass and, therefore, no ion dynamics), lattice distortions appear for infinitesimal e-ph coupling [14], as also previously found in the strong-coupling (Heisenberg) model [15]. A full quantum mechanical treatment of the model, including both electron and ion dynamics, has been addressed in a relatively limited number of works. In Ref. [16], it has been argued that an on-site electron repulsion suppresses the Peierls distortions for large phonon frequencies. Later, a few works, mainly using Monte Carlo or density-matrix renormalization group (DMRG) techniques, analyzed a variety of models with both quantum lattice fluctuations and short-range e-e interactions [17, 18, 19]. In particular, acoustic and optical phonons have been compared at half-filling (also considering different ways to couple lattice vibrations to the electronic hopping) [20]. The general outcome suggests that two phases can be stabilized: the Mott insulator (when the e-e interaction dominates), with no lattice distortions and gapless spin excitations, and the Peierls insulator (when the e-ph coupling dominates), characterized by lattice distortions, a two-fold degenerate ground state, and a fully gapped spectrum. The transition between them is described by the Kosterliz-Thouless universality class [21], making its precise location extremely challenging.
@@
Away from half-filling, a limited number of investigations have been attempted. Indeed, ground-state Monte Carlo techniques suffer from the sign problem, which prevents one from assessing large system sizes, and the DMRG accuracy is strongly affected by the large entanglement of the gapless (metallic) ground state. One interesting aspect is to understand whether a so-called Luther-Emery phase [22], namely a metal with gapped spin excitations, may ever emerge when doping the Peierls insulator. In this respect, analytic and numerical works demonstrated that the Luther-Emery liquid appears in the doped two-leg ladder Hubbard model (with no phonons) [23, 24, 25, 26, 27]. Its existence is triggered by the presence of multiple Fermi points in the non-interacting band structure, as discussed in weak-coupling approaches [23, 26]. For example, in the one-dimensional Hubbard model with both nearest- ($t$) and next-nearest-neighbor ($t^{\prime}$) hopping, a metallic phase with gapped spin excitations emerges by lightly doping the dimerized insulator found at $t^{\prime}>t/2$ [23, 24, 25]. In addition, Ref. [28] suggested that the Luther-Emery liquid may also emerge from doping an insulating phase with bond order obtained within a single-band Hubbard model in presence of both on-site and nearest-neighbor interactions. By analogy, one could expect that the same phase could appear upon doping the (dimerized) Peierls insulator stabilized by the electron-phonon interaction. An early work [29], based upon bosonization and renormalization-group methods, indeed showed that a spin-gapped metal emerges close to half-filling due to the phonon-assisted backward scattering. More recently, both static and dynamical properties have been addressed by DMRG [30]. However, a single electron density $n$ has been considered, relatively far away from half-filling (i.e., $n=0.75$), and Luttinger-liquid properties are observed, with gapless charge and spin degrees of freedom.
@@
In this work, we study the one-dimensional Hubbard-SSH model by means of a variational Monte Carlo approach (VMC). We employ a variational wave function composed of three terms: the first one is a simple phonon condensate, which can give rise to a staggered (dimerized) pattern of the lattice displacements; the second one is a Bardeen-Cooper-Schrieffer (BCS) state; the last one is a Jastrow factor, which includes electron-electron, phonon-phonon, and electron-phonon correlations. The main novelty of the variational Ansatz resides in the inclusion of “backflow” correlations within the BCS wave function, through an auxiliary BCS Hamiltonian which explicitly depends upon the phonon coordinates. We present a detailed analysis of the SSH model in presence of a local Hubbard interaction, both at half-filling and for various hole doping. At half-filling, the ground state is either a Mott or a Peierls insulator, depending on the values of the e-e and e-ph couplings, in agreement with previous calculations [17, 18, 19]. The central part of our work concerns the doped case, where metallic states appear. In fact, while doping a Mott insulator gives rise to a standard Luttinger liquid, similarly to what happens in the Hubbard model without phonons, in a lightly doped Peierls insulator the spin gap remains finite, thus stabilising a Luther-Emery liquid, which has never been noticed in this kind of coupled electron-phonon problems.
@@ II.2 The variational wave function
In particular, the sampling procedure includes both electron and phonon moves; within the former, either single or double (spin flip) electron hopping at nearest-neighbor sites are considered; the latter consists in local updates of the displacements $X_{j}\mapsto X_{j}+\Delta$, with $\Delta$ uniformly distributed within the interval $[-\Delta_{\rm max},\Delta_{\rm max}]$. The acceptance probability is then computed by the Metropolis algorithm.
Let us now discuss the specific construction of the variational wave function that is used in this work. In previous works [31, 32, 33], the electronic part has been taken as a pair-product state:
@@
where the hopping ($t_{ij}=t^{*}_{ji}$) and pairing ($\Delta_{ij}=\Delta_{ji}$) amplitudes are variational parameters, as well as the chemical potential $\mu$. Alternatively, it is possible to optimize directly the pairing function $f_{i,j}$, without passing through the definition of the auxiliary Hamiltonian [34]. Within these approaches, the electronic wave function does not depend upon the phonon displacements, e.g., the pairing function does not change along the Markov chain, once the variational parameters are kept fixed.
@@
where $g_{m}=g_{-m}$ and $h_{m}=-h_{-m}$ are variational parameters for $m=\pm 1$ and ±$\pm 3$ (suitable for a possible bond dimerization) and $\{X_{j}\}$ are the phonon displacements in the configuration $|X\rangle$ visited along the Markov chain. In this way, $\langle X;n_{\sigma}|\Psi_{\rm e}\rangle$ depends upon the actual phonon configuration. In this sense, a sort of “backflow” correlations are included in the electronic part of the variational wave function. In the standard case on the continuum, the effective position of every electron (from which the Slater determinant is constructed) depends on all the other ones [35, 36, 37]; on the lattice, this approach has been extended by constructing single-particle orbitals or pairing functions that explicitly depend upon the many-body electron configuration [38, 39]. Here, instead, backflow correlations involve electrons and phonons.
The phonon part has the same form as used in Ref. [33]:
@@
Finally, the e-e and e-ph correlations are included by standard Jastrow factors [34, 33]:
@@
where the pseudo-potentials $v_{ij}$, $u_{ij}$, and $w_{ij}$ are treated as variational parameters. It turns out that the optimal energy is obtained by imposing both translation and reflection symmetries, with $v_{ji}=v_{ij}$, $u_{ji}=u_{ij}$, and $w_{ji}=-w_{ij}$. The density-density Jastrow factor is particularly relevant to describe the Mott insulating phase, in which the Fourier transform of the pseudo-potential $v_{q}\approx 1/q^{2}$ for small momenta $q$ [40].
In the following, we consider clusters with $L$ sites and periodic-boundary conditions. All the variational parameters of the variational wave function are optimized by using the stochastic reconfiguration approach [41, 42]. Our variational wave function is implemented in a JAX-based code [43] that runs on parallel CPUs thanks to mpi4jax [44].
@@ III.1 Benchmarks with DMRG calculations
First, we would like to assess the accuracy of our approach and compare the variational energies to DMRG results at half-filling, for $L=50$. In the DMRG calculations, periodic-boundary conditions are taken on the Hamiltonian, but a truncation in the phonon Hilbert space is necessary. Here, we fix a maximum occupancy of $10$ bosons per site (which is always much larger than the average occupation). Then, the energy of the electron-phonon Hamiltonian is optimized by means of the DMRG algorithm implemented in the ITensor library [45] for a matrix-product state (MPS) which, for computational efficiency, has no translational symmetry. The accuracy of DMRG calculations is verified by evaluating the variance of the total energy, which is always below $0.007t^{2}$; the bond dimension of the MPS is always above $\chi=1200$, in some cases reaching $\chi=2400$.
@@ III.2 Results at half-filling
Let us now discuss the results at half-filling by focusing on the best variational Ansatz with backflow correlations. By varying the e-e repulsion $U$ and the e-ph interaction $\lambda$, the ground state is either a Mott (uniform) or a Peierls (dimerized) insulator [17]. Within our approach, the presence of a finite lattice distortion is signaled by the stabilization of a finite $z$ parameter in Eq. (13), which gives a dimerized pattern around which the phonon displacements are distributed. In this case, the translational symmetry is explicitly broken in the variational wave function and dimerization can be also detected from the calculation of the bond-order parameter:
@@
The results are summarized in the phase diagram shown in Fig. 2, fixing $\hbar\omega/t=1$. Here, the trivial limits are $\lambda=0$, which corresponds to the Mott insulator with no e-ph coupling, and $U=0$, which corresponds to the Peierls insulator. The transition between these two phases is of the Kosterliz-Thouless type [21]. Therefore, it is extremely difficult to locate its actual location within numerical calculations, since the spin gap of the Peierls phase is exponentially small close to the transition. This is particularly relevant in the vicinity of the non-interacting limit $U=\lambda=0$, where large clusters become necessary to detect the presence of the tiny dimerization that exists for $\lambda\to 0$. Nevertheless, sufficiently away from the transition line (and the non-interacting limit), our variational approach can straightforwardly distinguish two different regimes, with and without bond order.
@@
The results for $U/t=4$ are reported in Fig. 3 for two different regimes, i.e., $\lambda=0.04$ (Mott insulator) and $\lambda=0.25$ (Peierls insulator). The DMRG results (on the same cluster size, optimized for the Hamiltonian with periodic-boundary conditions) are also reported for comparison. First of all, the difference between Mott and Peierls states appears in the bond structure factor, which shows a huge peak at $q=\pi$ (i.e., $B(\pi)\approx 30$) in the latter case; the presence of a true dimerization implies a divergence of $B(\pi)$ with $L$. Within the variational approach, the huge peak at $q=\pi$ comes from the stabilization of a finite phonon parameter $z$ in Eq. (13), i.e. a finite lattice distortion. In other words, the existence of a Peierls phase can be more easily detected by looking at $z$ [or equivalently at $B_{\rm e}$ defined in Eq. (17)] than considering the bond structure factor $B(q)$. Within the Mott state, instead, $B(\pi)$ is much reduced, still having a cusp. Let us now turn to density and spin correlations. For both insulating phases, $N(q)$ is smooth and shows a quadratic behavior ($N(q)\propto q^{2}$) for small $q$, indicating that the charge excitations are gapped [40]. Also the spin structure factor $S(q)$ in the Peierls phase is smooth, signaling that also spin excitations are gapped. By contrast, in the Mott phase $S(q)$ is linear for small $q$ and has a peak at $q=\pi$, which is expected to diverge logarithmically with the system size and signals the presence of gapless spin excitations.